OpenStudy (anonymous):

7.4 - k = 5 ; k = 3.4 6th grade help

OpenStudy (anonymous):

Is it me or does it not equal 5?

OpenStudy (anonymous):

I was absent I missed when I was in school because I was sick. I don't know how do it there couple questions for my homework that I don't know how to do.

OpenStudy (anonymous):

don't be worry look at first equation you need to have k alone in one side first subtract both side of 5 then you will have a equation like this 3.4-k=0 now add both side with k 3.4-k+k=k you will reach 3.4=k

OpenStudy (mathstudent55):

The solution of the equation is k = 2.4 Why do you have k = 3.4?

OpenStudy (anonymous):

I'm confused a bit, does the problem actually say that k=3.4?

OpenStudy (anonymous):

Yea and idk

OpenStudy (anonymous):

yes 2.4-k=0 2.4-k+k=k k=2.4

OpenStudy (anonymous):

Do u want a picture of it?

OpenStudy (mathstudent55):

@Jay1274 Can you explain what you are trying to do? You have an equation and an incorrect solution. Is the problem simply the equation and k = 3.4 is what you got for k and you want to know if it's correct?

OpenStudy (anonymous):

It's solution of equations

OpenStudy (anonymous):

Are you trying to find if it's a solution or not, or are you trying to solve for k?

OpenStudy (anonymous):

Yea I think so

OpenStudy (mathstudent55):

Can you please write the instructions that came witht he problem?

OpenStudy (anonymous):

Okay but how can I put a picture on it I'm on my ipad

OpenStudy (anonymous):

Take a picture of it from your ipad and save it then click attach file on here.

OpenStudy (anonymous):

Where's the attach file

OpenStudy (mathstudent55):

If you are trying to find if k = 3.4 is the solution to the given equation, then substitute 3.4 for k in the equation. 7.4 - k = 5 7.4 - 3.4 = 5 4 = 5 This is false, so k = 3.4 is not a solution to the given equation.

OpenStudy (anonymous):

On the bottom you'll see three blue buttons on the left middle of the screen (equation, draw, and attach file). Click attach file.

OpenStudy (anonymous):

I don't see it

OpenStudy (mathstudent55):

If you are trying to solve the equation, then this is how you do it: 7.4 - k = 5 Subtract 7.4 from both sides: 7.4 - 7.4 - k = 5 - 7.4 -k = -2.4 Now multiply both sides by -1 k = 2.4 Solution is k = 2.4

OpenStudy (anonymous):

Ok @mathstudent55

OpenStudy (anonymous):

@LoveSpiderz

OpenStudy (anonymous):

Well, just go with what mathstudent55 said because he's right. Whatever the question is asking you. If you have to put in k for the equation then it's not a solution. If the question asked you find k just subtract 7.4 both sides (5) then you'll get k=2.4. Is there additional instructions to it like "Find if this is a solution" or "Solve for the following"?

OpenStudy (anonymous):

Okay @LoveSpiderz @mathstudent55

OpenStudy (anonymous):

Thanks for helping